# A Falsifiable Alternative to General Relativity

### v1

@@ 1 Introduction
Further motivation for modifying general relativity comes from the experimental sector. Astrophysical and cosmological observations indicate that retaining the standard gravitational action of general relativity requires the existence of undetected non-baryonic matter (dark matter) and an unknown form of energy (dark energy) comprising approximately $27\%$ and $68\%$ of the universe’s energy budget, respectively [5]. Moreover, various cosmological puzzles such as the horizon, flatness, and monopole problems indicate an early period of rapid cosmic expansion. One model that attempts to fit all of this data is $\Lambda$CDM (Lambda cold dark matter) supplemented by an inflationary scenario [6]. In addition to being plagued by the cosmological constant and coincidence problems, the $\Lambda$CDM model fails to explain the origin of the inflaton field or the nature of dark matter. Since $\Lambda$CDM is more of an ad-hoc empirical fit to astrophysical data, rather than a theoretically well-motivated theory, it has led many to consider modifying our description of gravity [7].
@@
where $\mathcal{R}$ is the Ricci scalar in the Palatini formalism given by $\mathcal{R}=g^{\mu\nu}\left(\partial_{\rho}\Gamma^{\rho}_{\nu\mu}-\partial_{\nu}\Gamma^{\rho}_{\rho\mu}+\Gamma^{\rho}_{\rho\lambda}\Gamma^{\lambda}_{\nu\mu}-\Gamma^{\rho}_{\nu\lambda}\Gamma^{\lambda}_{\rho\mu}\right)$ with the metric $g_{\mu\nu}$ and connection $\Gamma^{\sigma}_{\mu\nu}$ treated as independent variables [24]. The exponent $n(\mathcal{R})$ is a dimensionless function of $\mathcal{R}$ [20]. In the low-curvature limit (IR) AWIG recovers $f\left(\mathcal{R}\right)=\mathcal{R}$, which is identical to general relativity [6]. In the high curvature limit (UV) AWIG asymptotically approaches $f\left(\mathcal{R}\right)=\mathcal{R}^{2}$, which is a power-counting renormalizable and locally scale-invariant theory [20, 25]. AWIG is likely to be unitary because states of negative norm (ghosts) that cause unitarity violations do not appear in $f(R)$ theories [6, 26].
@@ 2 Completing the model
In $N-$dimensions, all locally scale-invariant $f(\mathcal{R})$ actions in the Palatini formalism take the form [29, 20]
@@
where the two parameters $a$ and $b$ are not necessarily equal. Expressing the action in this way quantifies the degree to which local scale invariance is broken via departures from the Weyl symmetric case $a=b$. A dimensionless action requires $[\kappa]=[\mathcal{R}^{a/2}]+[\sqrt{-g}]$, and since the canonical length dimension of $\mathcal{R}^{a/2}$ is $-a$ and for $\sqrt{-g}$ it is $b$, this implies $[\kappa]=b-a$ [^1] Here we have chosen the convention in which the metric tensor is assigned a dimensionality rather than the coordinates, which one is free to do since it is a matter of convention as to how to distribute the two canonical length dimensions of the line element $ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}$ across the constituent factors. Hence, the magnitude of the canonical length dimension of the gravitational coupling is a quantitative measure of the degree to which Weyl symmetry is broken.
The Einstein-Hilbert action of general relativity is a special case of Eq. (4), namely when $a=2$ and $b=4$ [6]. Since in this case $[\kappa]=2$, it is not locally scale-invariant. The fact that the gravitational coupling has a positive canonical length dimension can be viewed as the origin of the power-counting non-renormalizability of general relativity since it introduces a length scale into the theory. Making general relativity a locally scale-invariant theory at high energies therefore requires the canonical length dimension of the gravitational coupling to run from $[\kappa]=2$ in the IR to $[\kappa]=0$ in the UV. Since $a=2$ is fixed in general relativity, this scenario would require the spacetime dimension $b$ to run from $b=4$ in the IR to $b=2$ in the UV. Indeed, general relativity as a perturbative quantum field theory is known to be power-counting renormalizable and locally scale-invariant in two dimensions [30].
@@
Since the gravitational coupling has length dimensions $[\kappa]=b-a$, if we fix $a=2$ and allow $b$ to vary then $[\kappa]_{a}=b-2$, where $[\kappa]_{a}$ refers to the dimensionality of the gravitational coupling for a fixed exponent $a$. Conversely if we fix $b=4$ and allow $a$ to vary then $[\kappa]_{b}=4-a$. We conjecture that these two statements are mathematically equivalent. That is, $[\kappa]_{a}=[\kappa]_{b}$, which yields $a=6-b$. Identifying $a/2$ with the running exponent $n(R)$ initially defined in the metric formalism for convenience, we then find $2n\left(R\right)=6-b$, and hence
@@
In this way the Lagrangian density is expressed purely in terms of the Palatini scalar curvature, which conveys some numerical advantages particularly when it comes to exploring the models cosmological dynamics [25].
@@ 3 Curing the problems with Palatini gravity
We also highlight the fact that the Palatini formalism does not suffer from the so-called Dolgov-Kawasaki instability, regardless of the specific functional form of $f(\mathcal{R})$ [43]. In contrast, metric $f(R)$ gravity is susceptible to the Dolgov-Kawasaki instability, unless the specific condition $f^{\prime\prime}(R)\geq 0$ is met [43].
@@ 4 A tractable approximation
We begin by defining $\mathcal{R_{*}}$ as the dimensionless combination $\mathcal{R_{*}}\equiv c\mathcal{R}\equiv\mathcal{R}/\mathcal{R}_{0}$, with $\mathcal{R}_{0}$ a constant with canonical length dimensions of $[\mathcal{R}_{0}]=-2$ that represents the maximum value $\mathcal{R}$ can take. Thus, $\mathcal{R}_{*}$ is dimensionless and has a restricted domain of $[0,1]$. Moreover, by defining the dimensionless function
@@
Equating Eqs. (22) and (25) and solving for ϵ()italic-$\epsilon\left(\mathcal{R}_{*}\right)$ then yields
@@ 5 Testing AWIG using binary black hole mergers
$1\lesssim\left(\frac{\omega_{GR}}{\omega_{AWIG}}\right)_{max}\lesssim 1.41$ (32)
@@ 5.1.1 Hanford
t(s) $\omega_{GR}$ $\omega_{Han}$ $\omega_{GR}$/$\omega_{Han}$ $\sigma$ tension with $\omega_{GR}$
0.36000 $45.2\pm 10.7$ $44.6\pm 13.0$ $1.0\pm 0.4$ $0.0$
@@
Table 1: A table showing the 11 times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Hanford observatory $\omega_{Han}$ for GW150914. The ratio $\omega_{GR}$/$\omega_{Han}$ is given in column four and the tension between $\omega_{Han}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
@@ 5.1.2 Livingston
t(s) $\omega_{GR}$ $\omega_{Liv}$ $\omega_{GR}$/$\omega_{Liv}$ $\sigma$ tension with $\omega_{GR}$
0.36000 $44.1\pm 7.1$ $44.6\pm 14.1$ $1.0\pm 0.4$ $0.0$
0.37500 $51.7\pm 5.3$ $50.0\pm 16.3$ $1.0\pm 0.4$ $0.1$
@@
Table 2: A table showing the 11 times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Livingston observatory $\omega_{Liv}$. The ratio $\omega_{GR}$/$\omega_{Liv}$ is given in column four and the tension between $\omega_{Liv}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
@@ 5.2.1 Hanford
The gravitational wave strain $h(t)$ computed via full numerical relativity simulations for parameters specific to GW151226 and without any filtering is shown in Fig. 7a [51, 49] [^7] Unfortunately, simulations for GW151226 that include a band-pass filter were not available to the authors, although based on Appendix 7.2 the filter does not seem to have a significant impact anyhow. Here, GWOSC simulations without a band-pass filter are used for an observer at spatial infinity, at extrapolation order $N=4$ and using the dominant mode $l=2$, $m=2$ [51, 49]. The same numerical relativity data is used in Section 5.2.2.
@@
Results for our analysis of the Hanford observation of GW151226 are presented in Fig. 8 and Tab. 3. Figure 8 shows that the ratio $\omega_{GR}/\omega_{Han}$ for $t\leq-0.007s$ is consistent with the GR prediction of unity. For $-0.007s\leq t\leq-0.001s$ there is a statistically significant dip below unity, reaching as low as $0.8$. Close to the merge point ($t=0s$), there is a statistically significant rise above unity to a value of $1.3$. Again, this is consistent with the general behaviour predicted by AWIG.
@@
t(s) $\omega_{GR}$ $\omega_{Han}$ $\omega_{GR}$/$\omega_{Han}$ $\sigma$ tension with $\omega_{GR}$
@@
0.00000 $581.4\pm 77.6$ $460.0\pm 55.3$ $1.3\pm 0.2$ $1.3$
Table 3: A table showing the 11 times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Hanford observatory $\omega_{Han}$. The ratio $\omega_{GR}$/$\omega_{Han}$ is given in column four and the tension between $\omega_{Han}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
As can be seen in Tab. 3, at $t=0s$ there is a $1.3\ \sigma$ tension between $\omega_{Han}$ and $\omega_{GR}$, while there is only a $0.5\ \sigma$ tension between $\omega_{Han}$ and $\omega_{AWIG}$ at the same time value. At $t=-0.00441s$, there is a $1.9\ \sigma$ tension between $\omega_{Han}$ and $\omega_{GR}$, while at the same time value there is only a $0.5\ \sigma$ tension between $\omega_{Han}$ and $\omega_{AWIG}$. Therefore, these data points are once again more consistent with AWIG than GR.
@@ 5.2.2 Livingston
Results for our analysis of the Livingston observation of GW151226 are summarized in Fig. 9 and Tab. 4. Figure 9 shows that the ratio $\omega_{GR}/\omega_{Liv}$ for $t\leq-0.011s$ is consistent with the GR prediction of unity. For intermediate times $-0.011s\leq t\leq-0.001s$ there is a statistically significant dip below unity, consistently reaching as low as $0.7$. Close to the merge point ($t=0s$) there is a statistically significant rise above unity reaching a value of $1.2$. Again, this also matches the general behaviour AWIG predicts.
@@
t(s) $\omega_{GR}$ $\omega_{Liv}$ $\omega_{GR}$/$\omega_{Liv}$ $\sigma$ tension with $\omega_{GR}$
@@
0.00000 $582.0\pm 63.0$ $500.0\pm 28.7$ $1.2\pm 0.1$ $1.2$
Table 4: A table showing the 11 times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Livingston observatory $\omega_{Liv}$ for GW151226. The ratio $\omega_{GR}$/$\omega_{Liv}$ is given in column four and the tension between $\omega_{Liv}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
As can be seen in Tab. 4, at $t=0s$ there is a $1.2\ \sigma$ tension between $\omega_{Liv}$ and $\omega_{GR}$, however there is a slightly lower tension of $1.0\ \sigma$ between $\omega_{Liv}$ and $\omega_{AWIG}$ at the same point. At $t=-0.00985s$, there is a $2.7\ \sigma$ tension between $\omega_{Liv}$ and $\omega_{GR}$, while at the same point there is a $0.8\ \sigma$ tension between $\omega_{Liv}$ and $\omega_{AWIG}$. At $t=-0.00788s$, there is a $2.2\ \sigma$ tension with $\omega_{GR}$ but only a $0.5\ \sigma$ tension with $\omega_{AWIG}$. At $t=-0.00591s$, there is a $1.7\ \sigma$ tension with GR but only a $0.1\ \sigma$ tension with the AWIG prediction. Finally, at $t=-0.00197s$, there is a $2.1\ \sigma$ tension with GR but only a $0.4\ \sigma$ tension with the AWIG prediction. Therefore, all five data points favour AWIG over GR in this case.

### v2

@@ 1 Introduction
[…]
@@
where $\mathcal{R}$ is the Ricci scalar in the Palatini formalism given by $\mathcal{R}=g^{\mu\nu}\left(\partial_{\rho}\Gamma^{\rho}_{\nu\mu}-\partial_{\nu}\Gamma^{\rho}_{\rho\mu}+\Gamma^{\rho}_{\rho\lambda}\Gamma^{\lambda}_{\nu\mu}-\Gamma^{\rho}_{\nu\lambda}\Gamma^{\lambda}_{\rho\mu}\right)$ with the metric $g_{\mu\nu}$ and connection $\Gamma^{\sigma}_{\mu\nu}$ treated as independent variables [24]. The exponent $n(\mathcal{R})$ is a dimensionless function of $\mathcal{R}$ [20]. In the low-curvature limit (IR) AWIG recovers $f\left(\mathcal{R}\right)=\mathcal{R}$, which is identical to general relativity [6]. In the high curvature limit (UV) AWIG asymptotically approaches $f\left(\mathcal{R}\right)=\mathcal{R}^{2}$, which is a power-counting renormalizable and locally scale-invariant theory [20, 25]. [^1] Presently, we only claim that AWIG is superficially renormalizable as shown by simple power-counting arguments. Namely, since gravity as a perturbative quantum field theory in $d$-dimensional spacetime dictates that the momentum $p$ scales with loop-order $L$ via $\int p^{A-[G]L}dp$ [26], where $[G]$ is the canonical mass dimension of Newton’s coupling and $A$ is a process-dependent constant that is independent of $L$, then because $G]\to 0$ in the UV limit of AWIG the momenta $p$ no longer scales with $L$, making it at least power-counting renormalizable. AWIG is likely to be unitary because states of negative norm (ghosts) that cause unitarity violations do not appear in $f(R)$ theories [6, 27].
@@ 2 Completing the model
In NN-dimensions, all locally scale-invariant $f(\mathcal{R})$ actions in the Palatini formalism take the form [30, 20]
@@
where the two parameters $a$ and $b$ are not necessarily equal. Expressing the action in this way quantifies the degree to which local scale invariance is broken via departures from the Weyl symmetric case $a=b$. A dimensionless action requires $\kappa]=[\mathcal{R}^{a/2}]+[\sqrt{-g}]$, and since the canonical length dimension of $\mathcal{R}^{a/2}$ is a-a and for $\sqrt{-g}$ it is $b$, this implies $\kappa]=b-a$. [^2] Here we have chosen the convention in which the metric tensor is assigned a dimensionality rather than the coordinates, which one is free to do since it is a matter of convention as to how to distribute the two canonical length dimensions of the line element $ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}$ across the constituent factors. Hence, the magnitude of the canonical length dimension of the gravitational coupling is a quantitative measure of the degree to which Weyl symmetry is broken.
The Einstein-Hilbert action of general relativity is a special case of Eq. (4), namely when $a=2$ and $b=4$ [6]. Since in this case $\kappa]=2$, it is not locally scale-invariant. The fact that the gravitational coupling has a positive canonical length dimension can be viewed as the origin of the power-counting non-renormalizability of general relativity since it introduces a length scale into the theory. Making general relativity a locally scale-invariant theory at high energies therefore requires the canonical length dimension of the gravitational coupling to run from $\kappa]=2$ in the IR to $\kappa]=0$ in the UV. Since $a=2$ is fixed in general relativity, this scenario would require the spacetime dimension $b$ to run from $b=4$ in the IR to $b=2$ in the UV. Indeed, general relativity as a perturbative quantum field theory is known to be power-counting renormalizable and locally scale-invariant in two dimensions [31].
@@
Since the gravitational coupling has length dimensions $\kappa]=b-a$, if we fix $a=2$ and allow $b$ to vary then $\kappa]_{a}=b-2$, where $\kappa]_{a}$ refers to the dimensionality of the gravitational coupling for a fixed exponent $a$. Conversely if we fix $b=4$ and allow $a$ to vary then $\kappa]_{b}=4-a$. We conjecture that these two statements are mathematically equivalent. That is, $\kappa]_{a}=[\kappa]_{b}$, which yields a=6ba=6-b. Identifying $a/2$ with the running exponent $n(R)$ initially defined in the metric formalism for convenience, we then find $2n\left(R\right)=6-b$, and hence
@@
In this way the Lagrangian density is expressed purely in terms of the Palatini scalar curvature, which conveys some numerical advantages particularly when it comes to exploring the models cosmological dynamics [25]. [^4] Since $n\left(R\right)$ is independently defined via external physical principles, namely, dimensional reduction and Weyl invariance, and not by Eq. (14), the construction is logically non-circular.
@@ 3 Curing the problems with Palatini gravity
We also highlight the fact that the Palatini formalism does not suffer from the so-called Dolgov-Kawasaki instability, regardless of the specific functional form of $f(\mathcal{R})$ [44]. In contrast, metric $f(R)$ gravity is susceptible to the Dolgov-Kawasaki instability, unless the specific condition $f^{\prime\prime}(R)\geq 0$ is met [44]. Furthermore, considering AWIG as defined by Eq. (2), then at least for each fixed value of $n(\mathcal{R})$, we get a set of standard Palatini $\mathcal{R}^{n}$ theories, which are well-known to evade the Ostrogradsky instability problem due to their equivalence with scalar tensor theories that lead to two derivative equations of motion. Since the value of $n(\mathcal{R})$ is locally a constant at each spacetime point, this implies that at least locally AWIG evades the Ostragadski instability. However, it is important to note that globally, for non-constant $n(\mathcal{R})$, the presence of the Ostragadski instability in AWIG remains an open issue that needs further investigation.
@@ 4 A tractable approximation
We begin by defining $\mathcal{R_{*}}$ as the dimensionless combination $\mathcal{R_{*}}\equiv c\mathcal{R}\equiv\mathcal{R}/\mathcal{R}_{0}$, with $\mathcal{R}_{0}$ a constant with canonical length dimensions of $\mathcal{R}_{0}]=-2$ that represents the maximum value $\mathcal{R}$ can take. Thus, $\mathcal{R}_{*}$ is dimensionless and has a restricted domain of $[0,1]$. Moreover, by defining the dimensionless function
@@
Equating Eqs. (22) and (25) and solving for $\epsilon\left(\mathcal{R}_{*}\right)$ then yields
@@ 5 Testing AWIG using binary black hole mergers
$1\lesssim\left(\frac{\omega_{GR}}{\omega_{AWIG}}\right)_{max}\lesssim 1.41.$ (32)
@@ 5.1.1 Hanford
t(s) $\omega_{GR}$ $\omega_{Han}$ $\omega_{GR}$$\omega_{Han}$ $\sigma$ tension with $\omega_{GR}$
0.36000 $5.2\pm 10.7 44$$4.6\pm 13.0$ $1.0\pm 0.4$ $0.0$
@@
Table 1: A table showing the $1$ times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Hanford observatory $\omega_{Han}$ for GW150914. The ratio $\omega_{GR}$$\omega_{Han}$ is given in column four and the tension between $\omega_{Han}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
@@ 5.1.2 Livingston
t(s) $\omega_{GR}$ $\omega_{Liv}$ $\omega_{GR}$$\omega_{Liv}$ $\sigma$ tension with $\omega_{GR}$
0.36000 $4.1\pm 7.1 44$$4.6\pm 14.1$ $1.0\pm 0.4$ $0.0$
0.37500 $51.7\pm 5.3$ $50.0\pm 16.3$ $\pm 0.4 0.10.1$
@@
Table 2: A table showing the $1$ times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Livingston observatory $\omega_{Liv}$. The ratio $\omega_{GR}$$\omega_{Liv}$ is given in column four and the tension between $\omega_{Liv}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
@@ 5.2.1 Hanford
The gravitational wave strain $h(t)$ computed via full numerical relativity simulations for parameters specific to GW151226 and without any filtering is shown in Fig. 7a [52, 50] [^9] Unfortunately, simulations for GW151226 that include a band-pass filter were not available to the authors, although based on Appendix 7.2 the filter does not seem to have a significant impact anyhow. Here, GWOSC simulations [53] without a band-pass filter are used for an observer at spatial infinity, at extrapolation order $N=4$ and using the dominant mode $l=2$, $m=2$ [52, 50]. The same numerical relativity data is used in Section 5.2.2.
@@
Results for our analysis of the Hanford observation of GW151226 are presented in Fig. 8 and Tab. 3. Figure 8 shows that the ratio $\omega_{GR}/\omega_{Han}$ for $t\leq-0.007s$ is consistent with the GR prediction of unity. For $0.007s\leq t\leq-0.001s$ there is a statistically significant dip below unity, reaching as low as $0.8$. Close to the merge point ($t=0s$), there is a statistically significant rise above unity to a value of $1.3$. Again, this is consistent with the general behaviour predicted by AWIG.
@@
t(s) $\omega_{GR}$ $\omega_{Han}$ $\omega_{GR}$$\omega_{Han}$ $\sigma$ tension with $\omega_{GR}$
@@
0.00000 $581.4\pm 77.6$ $460.0\pm 55.3$ $\pm 0.2 1.31.3$
Table 3: A table showing the $1$ times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Hanford observatory $\omega_{Han}$. The ratio $\omega_{GR}$$\omega_{Han}$ is given in column four and the tension between $\omega_{Han}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
As can be seen in Tab. 3, at $t=0s$ there is a $1.3\ \sigma$ tension between $\omega_{Han}$ and $\omega_{GR}$, while there is only a $0.5\ \sigma$ tension between $\omega_{Han}$ and $\omega_{AWIG}$ at the same time value. At t=0.00441st=-0.00441s, there is a $1.9\ \sigma$ tension between $\omega_{Han}$ and $\omega_{GR}$, while at the same time value there is only a $0.5\ \sigma$ tension between $\omega_{Han}$ and $\omega_{AWIG}$. Therefore, these data points are once again more consistent with AWIG than GR.
@@ 5.2.2 Livingston
Results for our analysis of the Livingston observation of GW151226 are summarized in Fig. 9 and Tab. 4. Figure 9 shows that the ratio $\omega_{GR}/\omega_{Liv}$ for $t\leq-0.011s$ is consistent with the GR prediction of unity. For intermediate times $0.011s\leq t\leq-0.001s$ there is a statistically significant dip below unity, consistently reaching as low as $0.7$. Close to the merge point ($t=0s$) there is a statistically significant rise above unity reaching a value of $1.2$. Again, this also matches the general behaviour AWIG predicts.
@@
t(s) $\omega_{GR}$ $\omega_{Liv}$ $\omega_{GR}$$\omega_{Liv}$ $\sigma$ tension with $\omega_{GR}$
@@
0.00000 $582.0\pm 63.0$ $500.0\pm 28.7$ $\pm 0.1 1.21.2$
Table 4: A table showing the $1$ times at which the frequency predicted by general relativity $\omega_{GR}$ is compared with the frequency measured by the Livingston observatory $\omega_{Liv}$ for GW151226. The ratio $\omega_{GR}$$\omega_{Liv}$ is given in column four and the tension between $\omega_{Liv}$ and the general relativistic prediction $\omega_{GR}$ is given in column five.
As can be seen in Tab. 4, at $t=0s$ there is a $1.2\ \sigma$ tension between $\omega_{Liv}$ and $\omega_{GR}$, however there is a slightly lower tension of $1.0\ \sigma$ between $\omega_{Liv}$ and $\omega_{AWIG}$ at the same point. At t=0.00985st=-0.00985s, there is a $2.7\ \sigma$ tension between $\omega_{Liv}$ and $\omega_{GR}$, while at the same point there is a $0.8\ \sigma$ tension between $\omega_{Liv}$ and $\omega_{AWIG}$. At t=0.00788st=-0.00788s, there is a $2.2\ \sigma$ tension with $\omega_{GR}$ but only a $0.5\ \sigma$ tension with $\omega_{AWIG}$. At t=0.00591st=-0.00591s, there is a $1.7\ \sigma$ tension with GR but only a $0.1\ \sigma$ tension with the AWIG prediction. Finally, at t=0.00197st=-0.00197s, there is a $2.1\ \sigma$ tension with GR but only a $0.4\ \sigma$ tension with the AWIG prediction. Therefore, all five data points favour AWIG over GR in this case.